# Inverse scattering for repulsive potential and strong singular interactions

### v1

@@ Abstract
In a previous work of 2014 on a quantum system governed by the repulsive Hamiltonian, the author proved uniqueness for short-range interactions described by a scattering operator consisting of regular and singular parts. In this paper, the singular part is assumed to have much stronger singularities and the same uniqueness theorem is proved. By applying the time-dependent method invented by Enss and Weder in 1995, the high-velocity limit for a wider class of the scattering operator with stronger singularities also uniquely determines uniquely the interactions of a multi-dimensional system.
@@ 1 Introduction
The real-valued function $V^{\rm sing}\in\mathscr{V}^{\rm sing}$ is compactly supported in $\mathbb{R}^{n}$ and $V^{\rm sing}\in L^{q}(\mathbb{R}^{n})$, where
@@
$|\partial_{x}^{\beta}V^{\rm reg}(x)|\lesssim_{\beta}\langle x\rangle^{-|\beta|-\epsilon}$ (1.3)
@@
$W^{\pm}=\operatorname*{s-lim}_{t\rightarrow\pm\infty}e^{{\rm i}H}e^{-{\rm i}H_{0}}$ (1.4)
exist and are asymptotically complete by the result given in [4] (see also [17]), where $H=H_{0}+V$. Unfortunately, if $q=n/2$ for $n\geqslant 5$, $V^{\rm sing}$ is not always relatively $p^{2}$-compact but relatively $p^{2}$-bounded infinitesimally (see [23, Theorem X. 21]). Even in this instance, we prove in section 2 the existence of (1.4). We thus define the scattering operator such that
@@
Since the Enss–Weder time-dependent method was devised, many authors have applied it to establish the uniqueness of the interaction potentials for various quantum models. References [1], [2], [3], [11], [18], [19], [27], and [31] investigated models with external electric fields, whereas references [12] and [16] studied fractional and relativistic Laplacians, and [13] studied time-dependent harmonic oscillators. References [28], [29], and [30] applied the method to non-linear Schrödinger and Hartree–Fock equations. As for repulsive Hamiltonians, Theorem 1.2 was first proved in [20] assuming $V\in C^{\infty}(\mathbb{R}^{n})$ satisfies
@@ 2 High-dimensional case
In this section, we consider instances for which $n\geqslant 5$ and give a proof of Theorem 1.3. For this purpose, it suffices to prove the existence of the wave operators because we can demonstrate the proof of Theorem 1.3 in the same manner as given in [9]. We now start from the self-adjointness of $H$.
@@
$\leqslant\epsilon\|p^{2}({\rm i}+p^{2})^{-1}B_{1}({\rm i}+H_{0})\phi\|+C_{\epsilon}\|\chi({\rm i}+p^{2})^{-1}B_{1}({\rm i}+H_{0})\phi\|$
$\leqslant\epsilon\|B_{1}\|_{\mathscr{B}}\|H_{0}\phi\|+\epsilon\|B_{1}\|_{\mathscr{B}}\|\phi\|+C_{\epsilon}\|({\rm i}+H_{0})^{-1}B_{2}B_{1}({\rm i}+H_{0})\phi\|$ (2.2)
because $V^{\rm sing}$ is $p^{2}$-bounded infinitesimally. We find $({\rm i}+H_{0})^{-1}B_{2}B_{1}({\rm i}+H_{0})$ is bounded. This completes the proof. ∎
@@ 3 Low-dimensional case
Let $m\in H^{2}(\mathbb{R}^{n})$; that is, Sobolev space of order $2$. Then
$\sup_{\lambda\in\mathbb{R}}\|m(\lambda p)\|_{\infty}\lesssim\|m\|^{1-n/4}\|m\|_{\dot{H}^{2}}^{n/4}$ (3.2)
@@
$\|m(\lambda p)\|_{\infty}\lesssim\|m(\lambda x)\|^{1-n/4}\|m(\lambda x)\|_{\dot{H}^{2}}^{n/4},$ (3.3)
@@
by changing the variable $\eta=\xi/\lambda$. Noting that $(-n/2)(1-n/4)+(-n/2+2)(n/4)=0$, we have (3.2) from (3.3), (3.4) and (3.6). If $m\in H^{2}(\mathbb{R}^{n})$ generally, we take $m_{k}\in C_{0}^{\infty}(\mathbb{R}^{n})$ such that $m_{k}\rightarrow m$ as $k\rightarrow m$ in $H^{2}(\mathbb{R}^{n})$. Applying the limit in (3.2) for $m_{k}$, we have (3.2) for $m$ as in the proof of [21, Chap. 1, Theorem 3.1]. ∎
@@
by calculating the commutator $[\langle p\rangle^{-2},\chi(2x/\tau)]$. Noting that $V^{\rm sing}$ is compactly supported and that the right hand side of the first and second terms of (3.15) are equal to zero for $|v|\gg 1$, we have
@@
$\|\langle p/\sinh(2t)\rangle^{-2}\|_{\infty}\lesssim 1.$ (3.17)
@@
The Dollard-type modified wave operators were applied even for short-range inverse scattering under the electric fields by [18] and [19]. Later, the Graf-type [6] (or Zorbas-type [32]) modified wave operators were first introduced in [3] for the Stark effect instead of the Dollard-type. Since then, the Graf-type modification for inverse scattering under the electric fields has been the standard method (see also [1], [2] and [11]). Our modifier $e^{-{\rm i}\int_{0}^{t}V^{\rm reg}(\sinh(2\tau)v){\rm d}\tau}$ also commutes with any operator as with the Graf-type.
@@
$\|\langle(|v|/t)p\rangle^{-2}\|_{\infty}\lesssim 1.$ (3.41)
@@
performing the integration by parts $N$-times. The second terms is estimated and yeilds
@@
$\int_{|t|\geqslant|v|^{\sigma}}R_{2}{\rm d}t\lesssim|v|^{3+n/2-(2+N)\sigma}|\int_{|v|^{\sigma}}^{\infty}t^{-1-n/2}{\rm d}t=O(|v|^{3+n/2-(2+N+n/2)\sigma}).$ (3.47)

### v3

@@ Abstract
In a previous work of 2014 on a quantum system governed by the repulsive Hamiltonian, the author proved uniqueness for short-range interactions described by a scattering operator consisting of regular and singular parts. In this paper, the singular part is assumed to have much stronger singularities and the same uniqueness theorem is proved. By applying the time-dependent method invented by Enss and Weder in 1995, the high-velocity limit for a wider class of the scattering operator with stronger singularities also uniquely determines the interactions of a multi-dimensional system.
@@ 1 Introduction
The real-valued function $V^{\rm sing}\in\mathscr{V}^{\rm sing}$ is compactly supported in $\mathbb{R}^{n}$ and $V^{\rm sing}\in L^{q}(\mathbb{R}^{n})$ where
@@
$|\partial_{x}^{\beta}V^{\rm reg}(x)|\lesssim_{\beta}\langle x\rangle^{-\epsilon-|\beta|}$ (1.3)
@@
$W^{\pm}=\operatorname*{s-lim}_{t\rightarrow\pm\infty}e^{{\rm i}tH}e^{-{\rm i}tH_{0}}$ (1.4)
exist and are asymptotically complete by the result given in [4] (see also [17]), where $H=H_{0}+V$. If $q=n/2$ for $n\geqslant 5$, it is known that $V^{\rm sing}$ is relatively $p^{2}$-bounded infinitesimally (see [23, Theorem X. 21]). Although its relatively $p^{2}$-compactness also holds by applying Hardy-Littlewood-Sobolev inequality ([27, VIII 4.2]) even in this instance, we instead give the direct proof of the existence of (1.4) in section 2. We thus define the scattering operator such that
@@
Since the Enss–Weder time-dependent method was devised, many authors have applied it to establish the uniqueness of the interaction potentials for various quantum models. References [1], [2], [3], [11], [18], [19], [28], and [32] investigated models with external electric fields, whereas references [12] and [16] studied fractional and relativistic Laplacians, and [13] studied time-dependent harmonic oscillators. References [29], [30], [31], [33] and [34] applied the method to non-linear Schrödinger and Hartree–Fock equations. As for repulsive Hamiltonians, Theorem 1.2 was first proved in [20] assuming $V\in C^{\infty}(\mathbb{R}^{n})$ satisfies
@@ 2 High-dimensional case
In this section, we consider instances for which $n\geqslant 5$ and give a proof of Theorem 1.3. For this purpose, we give the direct proof of the existence of the wave operators as we mentioned before. We then can demonstrate the proof of Theorem 1.3 in the same manner as given in [9]. We now start from the self-adjointness of $H$.
@@
$\leqslant\epsilon\|p^{2}\chi({\rm i}+p^{2})^{-1}B_{1}({\rm i}+H_{0})\phi\|+C_{\epsilon}\|\chi({\rm i}+p^{2})^{-1}B_{1}({\rm i}+H_{0})\phi\|$
$=\epsilon\|p^{2}\chi({\rm i}+p^{2})^{-1}B_{1}({\rm i}+H_{0})\phi\|+C_{\epsilon}\|({\rm i}+H_{0})^{-1}B_{2}B_{1}({\rm i}+H_{0})\phi\|$ (2.2)
because $V^{\rm sing}$ is $p^{2}$-bounded infinitesimally. We find that $p^{2}\chi({\rm i}+p^{2})^{-1}B_{1}$ and $({\rm i}+H_{0})^{-1}B_{2}B_{1}({\rm i}+H_{0})$ are bounded. This completes the proof. ∎
@@ 3 Low-dimensional case
Let $m\in C^{\infty}(\mathbb{R}^{n})\cap H^{2}(\mathbb{R}^{n})$. Then
$\sup_{\lambda\in\mathbb{R}}\|m(\lambda p)\|_{\mathscr{B}(L^{\infty})}\lesssim\|m\|^{1-n/4}\|m\|_{\dot{H}^{2}}^{n/4}$ (3.2)
@@
$\|m(\lambda p)\|_{\mathscr{B}(L^{\infty})}\lesssim\|m(\lambda x)\|^{1-n/4}\|m(\lambda x)\|_{\dot{H}^{2}}^{n/4},$ (3.3)
@@
by changing the variable $\eta=\xi/\lambda$. Noting that $(-n/2)(1-n/4)+(-n/2+2)(n/4)=0$, we have (3.2) from (3.3), (3.4) and (3.6). If $m\in C^{\infty}(\mathbb{R}^{n})\cap H^{2}(\mathbb{R}^{n})$ generally, we take $m_{k}\in C_{0}^{\infty}(\mathbb{R}^{n})$ such that $m_{k}\rightarrow m$ as $k\rightarrow m$ in $H^{2}(\mathbb{R}^{n})$. Applying the limit in (3.2) for $m_{k}$, we have (3.2) for $m$ as in the proof of [21, Chap. 1, Theorem 3.1]. ∎
@@
by calculating the commutator $[\langle p\rangle^{-2},\chi(2x/\tau)]$. Noting that $V^{\rm sing}$ is compactly supported and that the first and second terms of the right-hand side of (3.15) are
$\int_{|v|\sinh 2}^{\infty}(\|V^{\rm sing}(x)\chi(2x/\tau)\langle p\rangle^{-2}\|_{\mathscr{B}}+\tau^{-1}\|V^{\rm sing}(x)(\nabla\chi)(2x/\tau)\langle p\rangle^{-2}\|_{\mathscr{B}}){\rm d}\tau=0$ (3.16)
for $|v|\gg 1$, we have
@@
$\|\langle p/\sinh(2t)\rangle^{-2}\|_{\mathscr{B}(L^{\infty})}\lesssim 1.$ (3.18)
@@
The Dollard-type modified wave operators were applied even for short-range inverse scattering under the electric fields by [18] and [19]. Later, the Graf-type [6] (or Zorbas-type [35]) modified wave operators were first introduced in [3] for inverse scattering in the Stark effect instead of the Dollard-type. Since then, the Graf-type modification for inverse scattering under the electric fields has been the standard method (see also [1], [2] and [11]). Our modifier $e^{-{\rm i}\int_{0}^{t}V^{\rm reg}(\sinh(2\tau)v){\rm d}\tau}$ also commutes with any operator as with the Graf-type.
@@
$\|\langle(|v|/t)p\rangle^{-2}\|_{\mathscr{B}(L^{\infty})}\lesssim 1.$ (3.42)
@@
performing the integration by parts $N$-times. The second terms is estimated and yields
@@
$\int_{|t|\geqslant|v|^{\sigma}}R_{2}{\rm d}t\lesssim|v|^{3+n/2-(2+N)\sigma}\int_{|v|^{\sigma}}^{\infty}t^{-1-n/2}{\rm d}t=O(|v|^{3+n/2-(2+N+n/2)\sigma}).$ (3.48)
@@
Data Availability. The data that support the findings of this study are available from the corresponding author upon reasonable request.